1. the triangle ABC is right-angled with right angle at C. the height from the corner C divides the hypotenuse into two parts with one length 1 and the other 4. What is the circumference of the triangle.
  2. given a right triangle with catheter lengths a and b, determine and enter the length of the bisector at the right angle.
  3. a rectangle has a diagonal length d. the acute angle between the diagonals is x. given that cos (x) = p, determine and state the length of the shorter sides of the rectangle as a function of p
  4. Given a > 1, what is the biggest solution to: (1/x-1) ≥ a/(x+1)
  5. Of cos x = p, and pi/2<x<pi. Give a expression for tan x
1
Expert's answer
2021-05-07T08:59:56-0400

1)


"a^2=h^2+1^2"

"b^2=h^2+4^2"

"5^2=a^2+b^2"

"5^2=1+b^2+16+b^2"

"2h^2=25-17=8"

"h^2=4"

"h=2"

so "a^2=2^2+1=5"

"a=2.236067977"


"b^2=2^2+16=20"

"b=4.472135955"


circumference

"=2.236067977+4.472135955+5=11.70820393"


2)




"BC=a, CA=b, AB=\sqrt{a^2+b^2}"


CD is the angle bisector

let's draw square CEDF and CE=h

so,

"BE=a-h , CD=h\sqrt{2}"


Now, ABC and DBE are similar triangles.

"\frac{b}{a}=\frac{h}{a-h}"


"h=\frac{ab}{a+b}"


length of angle bisector is given by

"CD=h\sqrt{2}"


"CD=\frac{ab\sqrt{2}}{a+b}"



3)




shortest sides of the rectangle are AB and DC

using trigonometric ratio in triangle ABC.

"<B=90\degree"

now,

"cosx=\frac{adjacent}{hypotenuse}"


"cosx=\frac{AB}{AC}"


"cosx=\frac{AB}{d}"


now, "p=\frac{AB}{d}"


"AB=pd"


length of shortest side "AB=p \times d"




4)

"a>1"


"(\frac{1}{x}≥\frac{a}{x+1}"


"\frac{1}{x}-x+\frac{1}{x}-1≥a"


"\frac{x}{x}-x+\frac{1}{x}-1≥a"


"1-x+\frac{1}{x}-1≥a"


"-x+\frac{1}{x}≥a"


"\frac{1}{x}-x≥a"


"\frac{1-x^2}{x}≥a"



5)

"cos x=p"


"\frac{1}{secx}=p"


"sec x=\frac{1}{p}"


"sec^2 x=(\frac{1}{p})^2"


By identity

"sec^2x-tan^2x=1"


"tan^2x=sec^2x-1"


"=\frac{1}{p^2}-1"


"=\frac{1-p^2}{p^2}"


"tan^2x=\frac{1-p^2}{p^2}"


"tanx =\sqrt{\frac{1-p^2}{p^2}}"


"tanx=\frac{\sqrt{1-p^2}}{p}"